OpenStudy (swaugh):

help i give medals Carrie and her friends hiked from the ranger's station due north for 3 miles then due west for 5 miles. If they hiked in a straight line back to the ranger's station, what distance would they travel? Round your answer to the nearest tenth of a mile.

OpenStudy (wolf1728):

Hello swaugh

OpenStudy (swaugh):

hi can you help me

OpenStudy (wolf1728):

Yes I can help you

OpenStudy (wolf1728):

They walk 3 miles north and then 5 miles west This forms a right triangle with legs of 5 and 3 The distance they would walk back to the ranger station is the hypotenuse of that triangle.

OpenStudy (swaugh):

these are the answer choices

OpenStudy (swaugh):

four point zero miles five point eight miles six point zero miles six point seven miles

OpenStudy (wolf1728):

hypotenuse^2 = 5^2 + 3^2 hypotenuse^2 = 25 + 9 hypotenuse^2 = 34 hypotenuse = sq root (34)

OpenStudy (wolf1728):

I think the person who made up that problem might have been a little confused. If they walked 3 miles north and then 4 miles west the hypotenuse would then equal 5 miles exactly.

OpenStudy (swaugh):

thats what i got and i dont know the right answer

OpenStudy (wolf1728):

it seems the answer would be square root of 34 or 5.8309518948 miles

OpenStudy (swaugh):

so 5.8

OpenStudy (wolf1728):

I guess 5.8 is close to the nearest tenth, so that's the answer.

OpenStudy (swaugh):

thank you

OpenStudy (wolf1728):

u r welcome